OpenStudy (anonymous):

Simplify by taking roots of the numerator and the denominator. Assume all expressions under radicals represent positive numbers. ^3sqrt8x^11/y^3

OpenStudy (anonymous):

or is the root not a real number

OpenStudy (anonymous):

\[\sqrt[3]({8x ^{11}}/y ^{3})=2x ^{11/3}/y=2x ^{3}\sqrt[3]{x ^{2}}/y\] hope u understood

OpenStudy (anonymous):

for the most part, yes :) Thank you